a student performing this experiment mistakenly used 6.0 ml of 16 M HNO3 to dissolve 0.18 g of solid copper, instead of 4.0ml described in the lab manual. what volume in milliliters of 6.0 M NaOH are required to neutralize the excess acid.
Answer #1


Balanced chemical equation is:
HNO3 + NaOH ---> NaNO3 + H2O


Here:
M(HNO3)=16 M
M(NaOH)=6.0 M
Volume of HNO3 to be neutralised by NaOH = excess volume
V(HNO3)=2.0 mL

According to balanced reaction:
1*number of mol of HNO3 =1*number of mol of NaOH
1*M(HNO3)*V(HNO3) =1*M(NaOH)*V(NaOH)
1*16.0 M *2.0 mL = 1*6.0M *V(NaOH)
V(NaOH) = 5.33 mL
Answer: 5.33 mL